OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–3, 7); Slope: 4

OpenStudy (anonymous):

y-y0=m(x-x0) we have y-7=4(x+3)

OpenStudy (anonymous):

then y=4x+19

OpenStudy (anonymous):

@kpopaigoo

OpenStudy (anonymous):

thank you @felavin